OpenStudy (coconut):

If f(x) = 12x + 7, what is f–1(x)?

Parth (parthkohli):

Do you mean f(-1)?

OpenStudy (unklerhaukus):

\[f^{-1}(x)\]

OpenStudy (anonymous):

lolol

OpenStudy (anonymous):

=(x-7)/12

OpenStudy (anonymous):

-f(x) maybe :P

OpenStudy (anonymous):

lmao f(x)^-1 :P

OpenStudy (anonymous):

or is that a negative prime? :o

Parth (parthkohli):

= \[\LARGE {1 \over 12x + 7}\]

Parth (parthkohli):

I'm not sure if right. If wrong, please correct me.

OpenStudy (anonymous):

icky

OpenStudy (anonymous):

that's wrong because the question is ambiguous

Parth (parthkohli):

Okay

OpenStudy (unklerhaukus):

inverse?

OpenStudy (anonymous):

ask coconut to set it up properly

OpenStudy (anonymous):

coconut is a coconut :P

Parth (parthkohli):

@Tai please don't speak of members like this. She/he will do it.

OpenStudy (coconut):

\[ f^{1} (x)\]i meant

OpenStudy (anonymous):

@ParthKohli OK

OpenStudy (anonymous):

12

OpenStudy (anonymous):

dy/dx = 12

OpenStudy (anonymous):

medal please

OpenStudy (anonymous):

12x^1 +7(x^0) --> 12 + 0

OpenStudy (anonymous):

f'(x) = 12

OpenStudy (anonymous):

medal please, one for me one for you :D

OpenStudy (coconut):

My options are x - 12/7 x+12/7 x-7/12 x+7/12

OpenStudy (unklerhaukus):

you haven't answered the question asked tai

OpenStudy (unklerhaukus):

\[f(x) = 12x + 7\]\[f(x)-7=12x\]\[\frac{f(x)-7}{12}=x\]\[\frac{x-7}{12}=f^{-1}(x)\]